OpenStudy (anonymous):

3(n+1)! divided by 5n!

OpenStudy (anonymous):

Please help, I need a good grade on my summer math packet.

OpenStudy (anonymous):

or this

OpenStudy (anonymous):

Do you think you could explain it, please?

OpenStudy (anonymous):

the answer is 3(n+1)/5

myininaya (myininaya):

\[\frac{3(n+1)!}{5n!}=\frac{3(n+1)n!}{5n!}=\frac{3(n+1)\not{n!}}{5\not{n!}}=\frac{3(n+1)}{5}\]

jimthompson5910 (jim_thompson5910):

(n+1)! = (n+1)(n)(n-1)...(3)(2)(1) n! = (n)(n-1)...(3)(2)(1) Notice how both have the terms (n)(n-1)...(3)(2)(1) in common, so they cancel.

jimthompson5910 (jim_thompson5910):

and they leave the term n+1 behind

jimthompson5910 (jim_thompson5910):

ie \[\frac{(n+1)!}{n!}=n+1\]

OpenStudy (anonymous):

\[\cos(\sin^{-1} 1/2)\]

OpenStudy (anonymous):

What about this one? I'm not sure where to start.

myininaya (myininaya):

make a right triangle

myininaya (myininaya):